OpenStudy (anonymous):

An element has three naturally occurring isotopes. Use the information below to calculate the weighted average atomic mass of the element, showing both the setup and the final answer for the calculation.

OpenStudy (cwrw238):

since i can see no information I'l illustrate the calculation in the case of Chlorine I'm relying on memory here so my figures may be a bit out. there are 2 naturally occurring isotopes of Chlorine with mass numbers 35 and 37. with the abundance of Cl 35 = 75% and of Cl 37 = 25% The average atomic mass is therefore [ 3(35) + 37] / 4 = 35.5

OpenStudy (aaronq):

in case your percentages of isotope abundance aren't good fractional values 1/3.. use this: % = percent abundance of isotope mass = mass of isotope (%) (mass 1) + (%) (mass 2) + (%) (mass 3) = Weighted average mass